OpenStudy (anonymous):

What is the distance between points (4, −7) and (−5, 9)? Round to the nearest tenth of a unit. units

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

what do you think it is?

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\) does this help ??

OpenStudy (anonymous):

i got 11.3 but it wont let me tpe the whole thing in

OpenStudy (anonymous):

noooo its kakarot

OpenStudy (anonymous):

and yea but what do i plug in?

hartnn (hartnn):

yeah I am both :)

hartnn (hartnn):

so you first find x1,x2,y1,y2 values 4,-7 means x1 = 4, y1= -7

hartnn (hartnn):

x2, y2 = -5,9 x2 = -5 y2 =9

OpenStudy (anonymous):

ohhhh

hartnn (hartnn):

4 - (-5) = ..? then square it -7-9 = ...? then square it

hartnn (hartnn):

lets complete cvliboi's question first :)

OpenStudy (anonymous):

9 and -16

hartnn (hartnn):

yep, now square them

OpenStudy (anonymous):

81 and 256

hartnn (hartnn):

add them and take the square root that will be it!

hartnn (hartnn):

you too :)

OpenStudy (anonymous):

c;

OpenStudy (anonymous):

18.3?

hartnn (hartnn):

very close, its 18.4 :)

OpenStudy (anonymous):

i cant put any decimals so just 18?

hartnn (hartnn):

thats right! :)

OpenStudy (anonymous):

thanks

hartnn (hartnn):

welcome ^_^